Coolbrook Company has the following information available for the past year:   

River Division Stream Division
Sales revenue $ 1,204,000 $ 1,807,000
Cost of goods sold and operating expenses 883,000 1,284,000
Net operating income $ 321,000 $ 523,000
Average invested assets $ 1,020,000 $ 1,460,000

   
The company’s hurdle rate is 6.26 percent.

Required:
2. Recalculate ROI and residual income for each division for each independent situation that follows: (Enter your ROI answers as a percentage rounded to two decimal places, (i.e., 0.1234 should be entered as 12.34%.))

a. Operating income increases by 8 percent.

River Division Stream Division
ROI % %
Residual Income (Loss)
Answer #1

2a) Calculate ROI and residual income (loss)

River Division Stream Division
ROI 346680/1020000 = 33.99% 564840/1460000 = 38.69%
Residual income 346680-(1020000*6.26%) = 282828 564840-(1460000*6.26%) = 473444